# Ghost projection via focal-field diffraction catastrophes

### v1

@@ V.4 Influence of Poisson noise
Any practical implementation of our ghost-projection technique would be susceptible to photon shot-noise, given by a Poisson distribution, due to quantization of radiant exposure. For the binary target images used earlier, the target signal sits at most around 30% above the pedestal in the reconstructions (e.g. panels (b), (d), and (f) of Fig. 7). For this case, letting $M$ represent the fraction of pixels that are non-zero, then from Eq. (18),
$\text{SNR}=\frac{(0.3)P\sqrt{M}}{\sqrt{1.3P}}\approx 0.26\sqrt{PM}.$ (19)
To exceed any desired SNR for these binary images, the projection needs to be exposed for long enough that $P$, the average number of photons per pixel in the pedestal, satisfies
$P\gtrapprox 14.4\frac{\text{SNR}^{2}}{M}.$ (20)
Figure 8 shows an example of this for multiple SNR values for the modified capital “R” target image. While we have here focused on the influence of SNR, for lithographic purposes, the required exposure may not be determined just by the SNR, but also the sensitivity of the lithographic film.
Figure 8: Simulated ghost projections of the modified capital “R” target image with Poisson noise, shown for four SNR values: (a) $\textrm{SNR}=0.1$, (b) $\textrm{SNR}=1$, (c) $\textrm{SNR}=3$, (d) $\textrm{SNR}=10$. The average number of photons $P$, incident on each pixel of pedestal, is displayed for each projection.
@@ VI Discussion
A statistical ensemble, say, of aberrated focal-field diffraction catastrophes, gives an overcomplete non-negative basis [^8] Here, “non-negative basis” refers to a set of non-negative pixellated intensity distributions that form a mathematical basis (modulo a pedestal) for pixellated intensity distributions when linear combinations of the set are formed with non-negative weighting coefficients. The stated linear combinations form non-negative intensity distributions, but when the pedestal is removed, the resulting distribution may be (and in general will be) non-negative. As an illustrative example, see Fig. 4(b) in Ref. [4], whose non-negative ghost-projection intensity distribution contains both negative and positive contributions relative to the pedestal. that may be used to synthesize any desired target intensity distribution. In this synthesis, the positive weighting coefficients $w_{k}$ are proportional to the exposure time for each pattern. The preceding two-sentence summary, of the core “ghost projection” [1, 2, 3, 4] idea that is explored in the present simulation-based paper, is qualified by the facts that (i) the synthesized target image has a spatial resolution that is governed by the finest features present in the diffraction-catastrophe ensemble, (ii) the target image is augmented by a smoothly-varying additive offset (or “pedestal”), and (iii) the transverse spatial extent of the synthesized image is limited by the transverse spatial extent of the basis patterns. Rather than seeking to shape the target field in a direct manner—as would be the case for image-synthesis using spatial light modulation, illumination of a single-patterned mask, or raster scanning of a tightly focused beam—our approach is indirect. In particular, our method recasts aberration-induced focal-field distortions as an enabling feature, when they can be created in a controllable and reproducible manner, since they generate the basis which underpins our indirect approach to focal-plane image synthesis. If we loosely speak of the aberration-catastrophe focal-field ensemble as a set of deterministic albeit highly structured “noise” maps, then our indirect approach may be spoken of as “building signals out of noise” [1] (cf. Fig. 2). Some particular approaches for achieving this means of beam shaping, for a variety of radiation and matter wave fields, are briefly explored in Sec. VI.1. Remarks on the efficiency of ghost projection using overcomplete basis sets are given in Sec. VI.2.
VI.1 Focal-field ghost projection using visible light, x rays, electrons and neutrons
Given that the focal-field ghost-projection method is applicable to a variety of radiation and matter wave fields, let us now comment on the associated differences in possible means for its future experimental implementation.
@@ VI.1 Focal-field ghost projection using visible light, x rays, electrons and neutrons
    In implementations for hard x rays, the phase modulator $B$ in Fig. 9 could consist of a transversely-displaced spatially-random amorphous screen that is weakly absorbing at such photon energies (e.g., a variable-thickness amorphous carbon film [32], a beryllium window [33] with smooth random variations in projected thickness, or a Kapton film [34] with variable projected thickness). An alternative means for phase modulation could employ a deformable mirror operating at grazing angles of incidence [35]. The focusing optic $C$ may be a Fresnel zone plate [36], compound refractive lens [37, 38], or Kirkpatrick–Baez mirror [39]. To obtain a sufficient degree of coherence, in all of the x-ray scenarios described above, energy-filtered x-ray synchrotron radiation could be employed.
@@
Figure 9: Coherent plane waves $A$ illuminate a smooth transparent spatially-random screen $B$, that may be reproducibly and accurately displaced in two directions transverse to the optical axis $z$. A focusing element $C$ then creates a diffraction catastrophe in the associated focal plane $D$. A suitable linear superposition of such focused fields yields a summed intensity distribution over the focal plane which approximates any desired target pattern, up to a limiting spatial resolution and additive pedestal.
VI.2 Efficiency of a diffraction-catastrophe basis
Complete orthogonal bases are often considered to be preferable to non-orthogonal bases, in a range of problems pertaining to synthesis or decomposition in the context of optical physics. This optimality of orthogonal complete bases may be challenged by nonorthogonal overcomplete [20] bases, when one wishes to synthesize one particular arbitrary function. In particular, one can choose a very small subset of a highly overcomplete basis, with the particular subset being chosen such that the representation of the ghost projection is sparse (i.e., the number of members in the subset is small). Thus stated, the ghost-projection method may be optimized via a reversed form of compressive sensing [6]. This reversed form of compressive sensing is implicitly performed in our NNLS approach to ghost projection, as evidenced by the large fraction of weights, $w_{k}$, that are set to zero by this approach (see Fig. 4(d)). Indeed, since the target image occupies a finite area $\mathcal{A}$ in the lens focal plane, there is a finite number of degrees of freedom
@@ VII Practical considerations and future work
Suppose that removing patterns with low weights (e.g. those below 1% of the maximum weight) has negligible effect on the synthesized image quality. The NNLS reconstructions assign large exposure times (weights) to a small number of patterns, so when the negligible patterns are omitted, each illuminating pattern is exposed for longer and the uncertainty introduced by the shutter speed is a smaller fraction of the exposure time for each pattern. This improves the resilience of the reconstruction to exposure-time noise. Additionally, if a non-negligible amount of time is required to adjust the phase modulation component between each illuminating pattern, then using fewer patterns decreases the required total projection time. It may be possible to reduce this adjustment time by ordering the illuminating patterns so that the phase changes minimally between consecutive patterns (i.e. the Zernike polynomial coefficients change minimally). Depending on the density of contributing illuminating patterns, there may need to be no beam downtime, and the illuminating pattern may be allowed to evolve continuously.
@@
Some applications, such as lithography, will require a spatially-uniform (“flat”) pedestal.
We present two possible avenues for pedestal flattening: (i) using the appropriate attenuating mask during projection, or (ii) superposing an inverted pedestal “after the fact”:
@@
    Assuming the pedestal to be independent of the target image, it should be possible to produce a single attenuating mask to insert between the lens and the focal plane such that the resulting projection has a flat pedestal. Using a material with attenuation coefficient $\mu$, the transmission through the mask is $e^{-\mu T_{1}(\mathbf{r})}$, where $T_{1}$ is the thickness of the mask along the direction of the optical axis which is perpendicular to the transverse coordinate $\mathbf{r}$. Let $\eta<\min\overline{I_{k}}(\mathbf{r})$ be the intensity of the desired uniform pedestal, i.e. we are looking for a thickness function, $T_{1}$, such that $\overline{I_{k}}(\mathbf{r})e^{-\mu T_{1}(\mathbf{r})}=\eta$. Rearranging, we get the necessary thickness of the mask
    $T_{1}(\mathbf{r})=\frac{1}{\mu}\ln\left(\frac{\overline{I_{k}}(\mathbf{r})}{\eta}\right).$ (22)
@@
    $\mathcal{P}^{*}(\mathbf{r})=C-\overline{I_{k}}(\mathbf{r})$ (23)
    where $C>\max\{\overline{I_{k}}\}$ is the value of the resulting flat pedestal. This can also be achieved using an attenuating mask. For a material with attenuation coefficient $\mu$ and transverse thickness function $T_{2}$, we require $\overline{I_{k}}(\mathbf{r})+Ae^{-\mu T_{2}(\mathbf{r})}=C$, where $A$ is the total uniform exposure incident upon the mask. The required thickness function is
    $T_{2}(\mathbf{r})=\frac{1}{\mu}\ln\left(\frac{A}{C-\overline{I_{k}}(\mathbf{r})}\right).$ (24)
    The advantage of this method is that as the pedestal flattening is separate from the image construction, the refractive properties of the attenuating mask will not warp the interference patterns, and so the precision in machining and placement of the mask does not need to be nearly as high. The disadvantage is that the resulting flat pedestal, $C$, is larger than the flat pedestal, $\eta$, in method (i), since $C>\max\overline{I_{k}}(\mathbf{r})>\min\overline{I_{k}}(\mathbf{r})>\eta$. This means that that the signal-to-noise ratio is preferable for method (i).
A number of avenues for future work remain. Experimental realization of the simulated work in this paper is necessary to establish the viability of using diffraction catastrophes for ghost-projection lithography. For instance, to compete with electron-beam lithography that has no pedestal, it must be demonstrated that the sub-Airy-disk detail within the diffraction catastrophes can be physically recreated to sufficient precision, as otherwise the fine details that make this method effective are not available for use. Furthermore, considerations such as the numerical aperture of the lens (or equivalent for other focusing systems) need to be incorporated. [^9] It is a standard optical-physics result that the resolution associated with a lens-based imaging system is on the order of the imaging-radiation wavelength $\lambda$ divided by the numerical aperture $N_{A}$ (see e.g. p. 144 of Ref. [17] or p. 466 of Ref. [11]). It would be interesting to seek a corresponding result for ghost projection via focal-field diffraction catastrophes, namely the relationship between $N_{A}$ and ghost-projection spatial resolution $s$. We conjecture that $s\approx F\lambda/N_{A}$, where $F$ is a dimensionless multiplicative factor on the order of unity, but further work is needed to investigate if this suggestion is correct. Also, the effect of imprecision in the phase modulation process and inaccuracy in the illuminating pattern exposure times requires investigation to determine usable parameters for practical purposes.
@@
Besides the mentioned applications in dynamic on-demand beam shaping, aberration correction, and lithography, a further potential application is tomographic volumetric additive manufacturing, which is an alternative technique to conventional layer-by-layer additive manufacturing (standard 3D printing) that involves solidifying a photopolymer by irradiating it from multiple angles with dynamic light patterns [46, 47, 48]. This would utilize the three-dimensional profile of the diffraction catastrophes, requiring analysis of how the diffraction patterns evolve with the distance along the beam axis. As demonstrated in e.g. Ref. [47], tomographic volumetric additive manufacturing already achieves smaller printing scales than layer-by-layer manufacturing, and if the material properties of the photopolymer are not the limiting factor, diffraction-catastrophe ghost projection may allow further miniaturization.
@@ VIII Conclusion
Ghost projection creates a desired distribution of radiant exposure by cumulatively exposing a fixed set of illuminating patterns. These basis-function patterns may have a random character, although the method may also employ nonrandom patterns. The technique developed in this study uses highly-structured diffraction catastrophes as basis functions, produced by focusing a coherent beam whose phase has been continuously modulated in a random manner. The approach, which may be viewed as reversed form of classical computational ghost imaging, has been demonstrated via simulation. Possible future applications of focal-field ghost projection include dynamic on-demand beam shaping of focused fields, an indirect form of aberration correction, and lithography.
@@ Acknowledgements.
AMK and DMP thank the Australian Research Council (ARC) for funding through the Discovery Project: DP210101312. AMK thanks the ARC and Industry partners funding the Industrial Transformation and Training Centre for Multiscale 3D Imaging, Modelling, and Manufacturing: IC180100008. This research utilized resources and services from the National Computational Infrastructure (NCI) facility, which is supported by the Australian Government. We acknowledge useful discussions with Christian Dwyer, Wilfred K. Fullagar, Kieran Larkin, Tanya M. Monro and Timothy C. Petersen.

### v2

@@ V.4 Influence of Poisson noise
Any practical implementation of our ghost-projection technique would be negatively influenced by photon shot-noise, given by a Poisson distribution, due to quantization of radiant exposure. For the binary target images used in Fig. 9, $\beta=0.1$ was used. Let $M$ denote the fraction of pixels in a binary target image that are white (with the remaining being black). If a white (i.e., maximum exposure) pixel of the target image receives an average of $K$ more photons than a black (i.e., minimum exposure) pixel of the target image at the same location, Eq. (20) implies
$\text{SNR}\approx\frac{K\sqrt{M}}{\sqrt{K(1+\frac{1-\beta}{\beta})}}=\sqrt{\beta KM}.$ (22)
To exceed any desired SNR for these binary images, the projection needs to be exposed for long enough that $K$ satisfies
$K\gtrapprox\frac{\text{SNR}^{2}}{\beta M}.$ (23)
Figure 10 shows an example of this for multiple $K$ values, using a Sierpiński carpet target image with $\beta=0.1$. While we have here focused on the influence of SNR, for lithographic purposes, the required exposure may be determined not only by the SNR but also by the sensitivity of the lithographic substrate.
Figure 10: Simulated ghost projections of a Sierpiński carpet target with $\beta=0.1$ in the presence of Poisson noise, shown for four K values: (a) $K=1000$, (b) $K=300$, (c) $K=100$, (d) $K=30$. The corresponding SNR values are displayed accordingly.
@@ VI Discussion
A statistical ensemble, say, of aberrated focal-field diffraction catastrophes, gives an overcomplete non-negative basis [^7] Here, “non-negative basis” refers to a set of non-negative pixellated intensity distributions that form a mathematical basis (modulo a pedestal) for pixellated intensity distributions when linear combinations of the set are formed with non-negative weighting coefficients. The stated linear combinations form non-negative intensity distributions, but when the pedestal is removed, the resulting distribution may be (and in general will be) both positive and negative. As an illustrative example, see Fig. 4(b) in Ref. [4], whose non-negative ghost-projection intensity distribution contains both negative and positive contributions relative to the pedestal. that may be used to synthesize any desired target intensity distribution. In this synthesis, the positive weighting coefficients $w_{k}$ are proportional to the exposure time for each pattern. The preceding two-sentence summary, of the core “ghost projection” [1, 2, 3, 4, 5] idea that is explored in the present simulation-based paper, is qualified by the facts that (i) the synthesized target image has a spatial resolution that is governed by the finest features present in the diffraction-catastrophe ensemble, (ii) the target image is augmented by a smoothly-varying additive offset (or “pedestal”), and (iii) the transverse spatial extent of the synthesized image is limited by the transverse spatial extent of the basis patterns. Rather than seeking to shape the target field in a direct manner—as would be the case for image synthesis using spatial light modulation, illumination of a single-patterned mask, or raster scanning of a tightly focused beam—our approach is indirect. In particular, our method recasts aberration-induced focal-field distortions as an enabling feature, when they can be created in a controllable and reproducible manner, since they generate the basis which underpins our indirect approach to focal-plane image synthesis. If we loosely speak of the aberration-catastrophe focal-field ensemble as a set of deterministic albeit highly structured “noise” maps, then our indirect approach may be spoken of as “building signals out of noise” [1] (cf. Fig. 2). Some particular approaches for achieving this means of beam shaping, for a variety of radiation and matter wave fields, are briefly explored in Sec. VI.1. Remarks on the advantages of a diffraction-catastrophe basis, together with the more general question of the efficiency of ghost projection using overcomplete basis sets, are given in Sec. VI.2.
VI.1 Focal-field ghost projection using visible light, x rays, neutrons, and electrons
Given that the focal-field ghost-projection method is applicable to a variety of radiation and matter wave fields, we now comment on associated differences in possible means for its future experimental implementation.
@@ VI.1 Focal-field ghost projection using visible light, x rays, neutrons, and electrons
    In implementations for hard x rays, the phase modulator $B$ in Fig. 11 could consist of a transversely-displaced spatially-random amorphous screen that is weakly absorbing at such photon energies (e.g., a variable-thickness amorphous carbon film [34], a beryllium window [35] with smooth random variations in projected thickness, or a Kapton film [36] with variable projected thickness). An alternative means for phase modulation could employ a deformable mirror operating at grazing angles of incidence [37]. The focusing optic $C$ may be a Fresnel zone plate [38], compound refractive lens [39, 40], or Kirkpatrick–Baez mirror [41]. To obtain a sufficient degree of coherence, in all of the x-ray scenarios described above, energy-filtered x-ray synchrotron radiation [42] could be employed.
@@
Figure 11: Coherent plane waves AA illuminate a smooth transparent spatially-random screen $B$, that may be reproducibly and accurately displaced in two directions transverse to the optical axis $z$. A focusing element $C$ then creates a diffraction catastrophe in the associated focal plane $D$. A suitable linear superposition of such focused fields yields a summed intensity distribution over the focal plane which approximates any desired target pattern, up to a limiting spatial resolution and additive pedestal.
VI.2 Advantages of a diffraction-catastrophe basis
Diffraction-catastrophe bases naturally arise from a beam-focusing configuration. First of all, they are straightforward to produce: any sufficiently strong perturbation of the beam’s phase will cause a highly-structured diffraction catastrophe at the focal plane, even if the phase perturbation itself is smooth and undetailed. Also, the patterns have been observed to contain features at a range of length scales, which is advantageous when constructing arbitrary images. It is difficult to produce some other basis (such as a binary speckle pattern) with such a setup. Physical attenuating masks are plagued by distorting diffraction effects at small scales (i.e. when the features are comparable to the wavelength). It is also impractical to reverse-engineer the necessary phase profile to project an arbitrary pattern upon focusing; not only is it a mathematically challenging task, but an arbitrary pattern will require an arbitrarily complicated phase profile, which may be impossible to accurately construct with the phase modulator. The diffraction catastrophe bases discussed in this paper require simple phase profiles by design, as the phase profiles are made using only the first 21 Zernike polynomials.
We now turn to a second point of advantage of the diffraction-catastrophe basis, which relates to its overcomplete nature. Complete orthogonal bases are often considered to be preferable to non-orthogonal bases, in a range of problems pertaining to synthesis or decomposition in the context of optical physics. This optimality of orthogonal complete bases may be challenged by nonorthogonal overcomplete [30] bases, when seeking to synthesize one particular arbitrary function. Indeed, one can choose a very small subset of a highly overcomplete basis, with the particular subset being chosen such that the representation of the ghost projection is sparse (i.e., the number of members in the subset is small). Thus stated, the ghost-projection method may be optimized via a reversed form of compressive sensing [14]. This reversed form of compressive sensing is implicitly performed in our NNLS approach to ghost projection, as evidenced by the large fraction of weights, $w_{k}$, that are set to zero by this approach (see Fig. 4(d)). Indeed, since the target image occupies a finite area $\mathcal{A}$ in the lens’ focal plane, there is a finite number of degrees of freedom
@@ VII Practical considerations and future work
Suppose that removing patterns with low weights (e.g. those below 1% of the maximum weight) has negligible effect on the synthesized image quality. The NNLS reconstructions assign large exposure times (weights) to a small number of patterns, so when the negligible-weight patterns are omitted, each illuminating pattern is exposed for longer and the uncertainty introduced by the shutter speed is a smaller fraction of the exposure time for each pattern. This improves the resilience of the reconstruction to exposure-time noise. Additionally, if a non-negligible amount of time is required to adjust the phase modulation component between each illuminating pattern, then using fewer patterns decreases the required total projection time. It may be possible to reduce this adjustment time by ordering the illuminating patterns so that the phase changes minimally between consecutive patterns (i.e. the Zernike polynomial coefficients change minimally). Depending on the density of contributing illuminating patterns, there may need to be no beam downtime, and the illuminating pattern may be allowed to evolve continuously.
@@
Some applications, such as lithography, will require a spatially-uniform (“flat”) pedestal. We present three possible avenues for pedestal flattening: (i) using an appropriate attenuating mask during projection, (ii) superposing an inverted pedestal “after the fact”, or (iii) sidestepping the issue by restricting the ghost projection field of view to be sufficiently small:
@@
    Assuming the pedestal to be independent of the target image, it should be possible to produce a single attenuating mask to insert between the lens and the focal plane such that the resulting projection has a flat pedestal. Using a material with attenuation coefficient $\mu$, the transmission through the mask is $e^{-\mu T_{1}(\mathbf{r})}$, where $T_{1}$ is the thickness of the mask along the direction of the optical axis which is perpendicular to the transverse coordinate $\mathbf{r}$. Let $\eta<\min\langle{I}(\mathbf{r})\rangle$ be the intensity of the desired uniform pedestal, i.e. we are looking for a thickness function, $T_{1}$, such that $\langle I(\mathbf{r})\rangle e^{-\mu T_{1}(\mathbf{r})}=\eta$. Rearranging, we get the necessary thickness of the mask
    $T_{1}(\mathbf{r})=\frac{1}{\mu}\ln\left(\frac{\langle{I}(\mathbf{r})\rangle}{\eta}\right).$ (25)
@@
    $\mathcal{P}^{*}(\mathbf{r})=C-\langle I(\mathbf{r})\rangle$ (26)
    where $C>\max\langle I(\mathbf{r})\rangle$ is the value of the resulting flat pedestal. This can also be achieved using an attenuating mask. For a material with attenuation coefficient $\mu$ and transverse thickness function $T_{2}$, we require $\langle{I}(\mathbf{r})\rangle+Ae^{-\mu T_{2}(\mathbf{r})}=C$, where $A$ is the total uniform exposure incident upon the mask. The required thickness function is
    $T_{2}(\mathbf{r})=\frac{1}{\mu}\ln\left(\frac{A}{C-\langle I(\mathbf{r})\rangle}\right).$ (27)
    The advantage of this method is that as the pedestal flattening is separate from the image construction, the refractive properties of the attenuating mask will not warp the interference patterns, and so the precision in machining and placement of the mask does not need to be nearly as high. The disadvantage is that the resulting flat pedestal, $C$, is larger than the flat pedestal, $\eta$, in method (i), since $C>\max\langle I(\mathbf{r})\rangle>\min\langle I(\mathbf{r})\rangle>\eta$. This means that that the signal-to-noise ratio is preferable for method (i).
  3. (iii)
    Pedestal flattening can also be achieved in a simple manner by confining the ghost projection to a region near the pedestal’s “center of mass” that has a diameter which is small compared to the width of the pedestal. In other words, the field of view can be cropped to a region over which the pedestal is approximately constant. This comes at the price of a lower resolution image.
A number of avenues for future work remain. Experimental realization of the simulated work in this paper is necessary to establish the viability of using diffraction catastrophes for ghost-projection lithography. For instance, to compete with electron-beam lithography that has no pedestal, it must be demonstrated that the sub-Airy-disk detail within the diffraction catastrophes can be physically recreated to sufficient precision, as otherwise the fine details that make this method effective are not available for use. Furthermore, considerations such as the numerical aperture of the lens (or equivalent for other focusing systems) need to be incorporated. [^8] It is a standard optical-physics result that the resolution associated with a lens-based imaging system is on the order of the imaging-radiation wavelength $\lambda$ divided by the numerical aperture $N_{A}$ (see e.g. p. 466 of Ref. [8] or p. 144 of Ref. [19]). It would be interesting to seek a corresponding result for ghost projection via focal-field diffraction catastrophes, namely the relationship between $N_{A}$ and ghost-projection spatial resolution $s$. We conjecture that $s\approx F\lambda/N_{A}$, where $F$ is a dimensionless multiplicative factor on the order of unity, but further work is needed to investigate if this suggestion is correct. Also, the effect of imprecision in the phase modulation process and inaccuracy in the illuminating pattern exposure times requires investigation to determine usable parameters for practical purposes.
@@
Besides the mentioned applications in dynamic on-demand beam shaping, aberration correction, and lithography, a further potential application is tomographic volumetric additive manufacturing, which is an alternative technique to conventional layer-by-layer additive manufacturing (standard 3D printing) that involves solidifying a photopolymer by irradiating it from multiple angles with dynamic light patterns [49, 50, 51]. This would utilize the three-dimensional profile of the diffraction catastrophes, requiring analysis of how the diffraction patterns evolve with distance along the beam axis. As demonstrated in e.g. Ref. [50], tomographic volumetric additive manufacturing already achieves smaller printing scales than layer-by-layer manufacturing, and if the material properties of the photopolymer are not the limiting factor, diffraction-catastrophe ghost projection may allow further miniaturization.
@@ VIII Conclusion
Ghost projection creates a desired distribution of radiant exposure by cumulatively exposing a fixed set of illuminating patterns. These basis-function patterns may have a random character, although the method may also employ nonrandom patterns. The technique developed in this study uses highly-structured diffraction catastrophes as basis functions, produced by focusing a coherent beam whose phase has been continuously modulated in a random manner. The approach, which may be viewed as a reversed form of classical computational ghost imaging, has been demonstrated via simulation. Possible future applications of focal-field ghost projection include dynamic on-demand beam shaping of focused fields, an indirect form of aberration correction, and lithography.
@@ Acknowledgements.
AMK and DMP thank the Australian Research Council (ARC) for funding through the Discovery Project: DP210101312. AMK thanks the ARC and Industry partners funding the Industrial Transformation and Training Centre for Multiscale 3D Imaging, Modelling, and Manufacturing: IC180100008. This research utilized resources and services from the National Computational Infrastructure (NCI) facility, which is supported by the Australian Government. We acknowledge useful discussions with Christian Dwyer, Wilfred K. Fullagar, Kieran Larkin, Tanya M. Monro, and Timothy C. Petersen.